# A Systematic Literature Review on Large Language Models for Automated Program Repair

### v3

@@ 1. Introduction
Very recently, Large Language Models (LLMs) have been successfully applied to a broad range of source code-related tasks (Zhang et al., 2023c; Wang et al., 2024a; Zhang et al., 2023a), such as code generation (Li et al., 2023b; Wang et al., 2023c; Zhu et al., 2024; Wang et al., 2023b), code summarization (Su and McMillan, 2024; Wang et al., 2023a; Sun et al., 2023), and test generation (Schafer et al., 2024; Dinella et al., 2022; Nashid et al., 2023; Hu et al., 2023; Alagarsamy et al., 2023). Benefiting from massive model parameters and vast training data, LLMs have demonstrated impressive performance and fundamentally revolutionized the research paradigm in the Software Engineering (SE) community. In the domain of APR, beginning with pioneering studies, e.g., TFix (Berabi et al., 2021), CIRCLE (Yuan et al., 2022) and AlphaRepair (Xia and Zhang, 2022), the community has witnessed an explosion of repair studies utilizing LLMs, already achieving considerable advantages and further indicating significant potential for future research. However, the integration of LLMs within APR is a considerably complex undertaking, making it difficult for interested researchers to understand existing work. For example, existing LLM-based APR studies encompass different research perspectives (e.g., empirical (Xia et al., 2023b), technical (Xia and Zhang, 2022) and benchmark studies (Zhang et al., 2023k)), repair phases (e.g., patch generation (Zhang et al., 2023e) and correctness assessment (Zhang et al., 2024b)), repair scenarios (e.g., static warnings (Jin et al., 2023) and syntax errors (Joshi et al., 2023)), mode architectures (e.g., encoder-only (Zhang et al., 2023d) and decoder-only (Mashhadi and Hemmati, 2021)) and model utilization paradigms (e.g., fine-tuning (Yuan et al., 2022), few-shot (Nashid et al., 2023) and zero-shot (Zhang et al., 2023e)). Despite ongoing explorations in the field, the literature currently lacks a detailed and systematic review of the applications of LLMs in APR, making it challenging for researchers to understand the multitudinous design choices of existing work and conduct follow-up research.
@@ 2.2. Large Language Models
LLMs refer to advanced Artificial Intelligence (AI) models that undergo extensive pre-training on large text corpora to enhance their capabilities of natural language understanding, generation, and interpretation in a human-like manner (Zhao et al., 2023b). In recent years, their training scope has been extended beyond natural languages to include programming languages, enabling LLMs to reason about and generate source code effectively. The literature has seen a variety of LLMs supporting NLP and SE research, which can be categorized into three main categories based on their model architectures: (1) Encoder-only LLMs, such as CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), train the encoder part of the Transformer to generate a fixed-dimensional bidirectional representation with Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). MLM aims to predict the original tokens that have been randomly masked out, and NSP predicts whether two given sentences actually follow each other in a text. (2) Decoder-only LLMs, such as CodeGPT (Lu et al., 2021), train the decoder part of the Transformer to support auto-regressive tasks with Causal Language Modeling (CLM), which aims to predict new tokens in a sequence based on previous tokens. (3) Encoder-decoder LLMs, such as CodeT5 (Wang et al., 2021), train both encoder and decoder parts of the Transformer to support sequence-to-sequence generation tasks with denoising objectives. We will summarize existing LLMs and how they are leveraged to support program repair in Section 6.1.
@@ 4.2. Search Strategy
We search and identify primary papers following the “Quasi-Gold Standard” (QGS) (Zhang et al., 2011) method, which is a common practice to construct a set of known studies for refining search strings by combining manual and automated search processes. We first conduct a manual search to identify a set of relevant studies and derive a search string from them, as detailed in Section 4.3. We then utilize the search string to perform an automated search and employ a series of relatively strict filtering steps to obtain the most relevant studies, as detailed in Section 4.4. We finally use a snowballing search to supplement the search results further, as detailed in Section 4.5. Given the abundance of relevant papers from different research communities, such as SE and AI, this search strategy allows us to capture the most pertinent papers while achieving greater efficiency than a purely manual search with a relatively rigorous process. Particularly, we undertake the following five phases to search for and identify relevant studies.
@@ 4.3. Search Items
To perform the manual search and establish QGS, we select four top-tier SE conferences (ICSE, ESEC/FSE, ASE, ISSTA) and two journals (TOSEM and TSE) and search for relevant papers involving both program repair and LLMs. We manually inspect and identify 25 papers from 2020 to 2024 that meet our criteria, which then form the basis for constructing QGS. Following prior surveys (Wang et al., 2022; Yang et al., 2022; Zhang et al., 2023b), we divide the search items used for searching papers automatically into two groups: (1) an APR-related group containing some commonly-used keywords related to program repair; and (2) an LLM-related group containing some popular keywords related to large language models or pre-trained models. Besides, as the APR community benefits from well-maintained forums and real-time updated works, we further refine our search items by concluding some frequent keywords from three sources: a community-driven website [^2] http://program-repair.org/bibliography.html, a living review of APR by Monperrus (Monperrus, 2022) and the most recent learning-based APR survey (Zhang et al., 2023b). Finally, we identify a search string that includes several LLM-related and APR-related keywords frequently appearing in program repair studies that use LLMs. The complete set of search keywords is as follows:
@@ 4.4.2. Quality Assessment
To ensure the inclusion of high-quality papers, we conduct a full-text inspection of the remaining papers to assess their relevance and quality within the LLM-based APR field. Following Wang et al. (Wang et al., 2022), we design ten quality assessment questions, where each paper is independently rated by two authors on a three-tier scale:“yes” (1 point), “partial” (0.5 points), or “no” (0 points). Papers scoring below the threshold of 8 points are excluded from further analysis, resulting in 167 papers being retained. Any disagreement is resolved through discussion, and when necessary, by consultation with the third author. The designed quality assessment criteria (QAC) are presented in Table 2. It is worth noting that, for QAC4, given the emerging nature of this research area and the rapid evolution of LLMs, many recent works have not yet completed the peer-review process. Therefore, we also consider papers from arXiv that satisfy our quality assessment criteria. Before inclusion, we carefully check whether each arXiv paper had been subsequently published by comparing its title, authorship, and content. If a published version is found, only the peer-reviewed version is retained; otherwise, the arXiv version is included and explicitly marked as such in our dataset. This strategy aligns with the common practice in recent LLM-based SE surveys, and ensures our work remains both comprehensive and up to date while maintaining a high standard of quality.
@@ 4.5. Snowballing Search
To ensure maximal comprehensiveness of the collected papers, we additionally employ the snowballing search  (Watson et al., 2022) to manually incorporate any potential papers that are overlooked during our previous search process. The snowballing search is a common practice to examine the references (i.e., backward snowballing) or citations (i.e., forward snowballing) within a list of papers to discover additional relevant works. Specifically, we scrutinize every reference and citation within the collected papers to assess their relevance to our research objectives. This meticulous manual analysis of all potential papers through the entire study selection process ultimately leads to the inclusion of 189 papers in our survey. All included papers and their bibliographic information are publicly available in our replication repository for transparency and reproducibility (Author, s).
@@ 5.4. What are the types of publication contributions?
We categorize collected papers according to their main contributions into four categories: new technique or methodology, empirical study, benchmark, and human study, as illustrated in Table 4. It is worth noting that some studies may fall into multiple categories, e.g., papers that first conduct an empirical evaluation and propose a new repair framework. We find that 118 relevant papers are published, with the aim of proposing a novel repair approach or framework with LLMs to address various issues in the APR community. Besides, 56 papers concentrate on conducting empirical studies to explore the actual benefits of LLMs in fixing various bugs, such as the potential of fine-tuning LLMs in vulnerability repair (Zhang et al., 2023d; Ge and Wu, 2023). We further notice 15 relevant studies constructing new benchmarks to evaluate the performance of LLMs and three papers (Ge and Wu, 2023; Meem et al., 2024) administering a survey to offer insights into how practitioners or developers think about and employ LLMs to fix software bugs in practice.
@@ 6.2. What approaches are employed to optimize LLMs for program repair?
2025 6 (Ehsani et al., 2025b), (Liu et al., 2025), (Orvalho et al., 2025), (Nong et al., 2025), (Lin et al., 2025), (Zhang et al., 2025)
@@
2025 17 (Deligiannis et al., 2023), (Garg et al., 2023), (Huang et al., 2025c), (Mündler et al., 2024), (Charalambous et al., 2023), (Xu et al., 2023), (Xia et al., 2025), (Yuan et al., 2024), (Zhang et al., 2024c), (Steenhoek et al., 2025), (Cao et al., 2023), (Hu et al., 2025), (Kong et al., 2024), (Li et al., 2025), (Ouyang et al., 2025), (Parasaram et al., 2025), (Kang et al., 2023)
@@ 6.2.4. Agent
The second category involves repository-level repair agents, which target real-world software issues, guided by the development of new benchmarks, such as SWE-bench (Jimenez et al., 2024). Unlike traditional approaches (Zhang et al., 2023b) that often rely on predefined conditions (e.g., assuming known bug locations, failure-triggering tests, or single-function repair assumptions), these agents require end-to-end bug repair across full repositories, requiring reasoning across multiple files, dependencies, and test suites. Recent studies have proposed a series of SWE-oriented techniques to address the challenges of repository-level repair, including designing specialized agent-computer interfaces (e.g., SWE-Agent (Yang et al., 2024a)), encoding the whole repository via AST representations (e.g., AutoCodeRover (Zhang et al., 2024d)), enabling multi-agent collaboration (e.g., MAGIS (Tao et al., 2024)), and developing generalist and specialist agents (e.g., OpenHands (Wang et al., 2024b)).
@@ 7. RQ3: What repair scenarios have been facilitated by LLMs?
2025 22 (Bouzenia et al., 2024) (Cao et al., 2023) (Ehsani et al., 2025a) (Feng et al., 2025) (Hu et al., 2025) (Kong et al., 2024) (Kong et al., 2025) (Lee et al., 2024) (Li et al., 2025) (Lin et al., 2025) (Luo et al., 2024) (Macháček et al., 2025) (Ouyang et al., 2025) (Parasaram et al., 2025) (Ruiz et al., 2025) (Xu et al., 2025) (Yang et al., 2024e) (Ye et al., 2025) (Huang et al., 2025b) (Zhang et al., 2025) (Kang et al., 2023) (Silva et al., 2023)
@@
2025 11 (Ehsani et al., 2025b) (Guo et al., 2025) (Ma et al., 2025a) (Mündler et al., 2024) (Antoniades et al., 2024) (Ouyang et al., 2024b) (Wei et al., 2025) (Xia et al., 2025) (Tao et al., 2024) (Wang et al., 2024b) (Ma et al., 2025b)
@@ 7.1. Semantic Bugs
2025 ReinFix (Zhang et al., 2025) ICSE GPT-3.5,GPT-4o,GPT-4 Java URL
@@ 7.4. Repository-level Issues
2025 Agentless (Xia et al., 2025) FSE/ESEC Approach GPT-4o Python URL
@@
Repository-level issues refer to real-world issues (such as bug reports or feature requests) in large-scale repositories. Recent LLM-based APR research increasingly targets repository-level issues, whose resolution requires long-context understanding over a large codebase, accurate fault localization from informal reports, dependency, and code edits across multiple modules, and robust validation of the resulting pull request. A prominent benchmark for this scenario is SWE-bench (Jimenez et al., 2024), which comprises 2294 real GitHub issues with paired ground-truth pull requests from 12 popular Python repositories. In addition, several other datasets, such as SWT-bench (Mündler et al., 2024), OmniGIRL (Guo et al., 2025), and SWE-bench M (Yang et al., 2024b), are used to evaluate LLMs under varying APR settings. To address repository-level repair challenges, researchers have explored two main categories of approaches: procedural (Xia et al., 2025; Ouyang et al., 2024b) and agentic (Zhang et al., 2024d; Antoniades et al., 2024; Chen et al., 2024b; Wang et al., 2024b). In particular, similar to few-shot or zero-shot prompting approaches, procedural approaches rely on expert-designed workflows where LLMs run at a fixed pipeline, and every choice is hard-coded. Agentic approaches query LLMs to autonomously plan and execute each choice by equipping LLMs with various external tools

### v4

@@ 1. Introduction
Very recently, Large Language Models (LLMs) have been successfully applied to a broad range of source code-related tasks (Zhang et al., 2023c; Wang et al., 2024a; Zhang et al., 2023a), such as code generation (Li et al., 2023b; Wang et al., 2023c; Zhu et al., 2024; Wang et al., 2023b), code summarization (Su and McMillan, 2024; Wang et al., 2023a; Sun et al., 2023), and test generation (Schafer et al., 2024; Dinella et al., 2022; Nashid et al., 2023; Hu et al., 2023; Alagarsamy et al., 2023; Zhang et al., 2025e, c, b). Benefiting from massive model parameters and vast training data, LLMs have demonstrated impressive performance and fundamentally revolutionized the research paradigm in the Software Engineering (SE) community. In the domain of APR, beginning with pioneering studies, e.g., TFix (Berabi et al., 2021), CIRCLE (Yuan et al., 2022) and AlphaRepair (Xia and Zhang, 2022), the community has witnessed an explosion of repair studies utilizing LLMs, already achieving considerable advantages and further indicating significant potential for future research. However, the integration of LLMs within APR is a considerably complex undertaking, making it difficult for interested researchers to understand existing work. For example, existing LLM-based APR studies encompass different research perspectives (e.g., empirical (Xia et al., 2023b), technical (Xia and Zhang, 2022) and benchmark studies (Zhang et al., 2023k)), repair phases (e.g., patch generation (Zhang et al., 2023e) and correctness assessment (Zhang et al., 2024b)), repair scenarios (e.g., static warnings (Jin et al., 2023) and syntax errors (Joshi et al., 2023)), mode architectures (e.g., encoder-only (Zhang et al., 2023d) and decoder-only (Mashhadi and Hemmati, 2021)) and model utilization paradigms (e.g., fine-tuning (Yuan et al., 2022), few-shot (Nashid et al., 2023) and zero-shot (Zhang et al., 2023e)). Despite ongoing explorations in the field, the literature currently lacks a detailed and systematic review of the applications of LLMs in APR, making it challenging for researchers to understand the multitudinous design choices of existing work and conduct follow-up research.
@@ 2.2. Large Language Models
LLMs refer to advanced Artificial Intelligence (AI) models that undergo extensive pre-training on large text corpora to enhance their capabilities of natural language understanding, generation, and interpretation in a human-like manner (Zhao et al., 2023b). In recent years, their training scope has been extended beyond natural languages to include programming languages, enabling LLMs to reason about and generate source code effectively. The literature has seen a variety of LLMs supporting NLP and SE research (Shang et al., 2025), which can be categorized into three main categories based on their model architectures: (1) Encoder-only LLMs, such as CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), train the encoder part of the Transformer to generate a fixed-dimensional bidirectional representation with Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). MLM aims to predict the original tokens that have been randomly masked out, and NSP predicts whether two given sentences actually follow each other in a text. (2) Decoder-only LLMs, such as CodeGPT (Lu et al., 2021), train the decoder part of the Transformer to support auto-regressive tasks with Causal Language Modeling (CLM), which aims to predict new tokens in a sequence based on previous tokens. (3) Encoder-decoder LLMs, such as CodeT5 (Wang et al., 2021), train both encoder and decoder parts of the Transformer to support sequence-to-sequence generation tasks with denoising objectives. We will summarize existing LLMs and how they are leveraged to support program repair in Section 6.1.
@@ 4.2. Search Strategy
Figure 2. Process of paper collection and selection.
We search and identify primary papers following the “Quasi-Gold Standard” (QGS) (Zhang et al., 2011) method, which is a common practice to construct a set of known studies for refining search strings by combining manual and automated search processes. Fig. 2 illustrates the general workflow of paper collection and section. We first conduct a manual search to identify a set of relevant studies and derive a search string from them, as detailed in Section 4.3. We then utilize the search string to perform an automated search and employ a series of relatively strict filtering steps to obtain the most relevant studies, as detailed in Section 4.4. We finally use a snowballing search to supplement the search results further, as detailed in Section 4.5. Given the abundance of relevant papers from different research communities, such as SE and AI, this search strategy allows us to capture the most pertinent papers while achieving greater efficiency than a purely manual search with a relatively rigorous process. Particularly, we undertake the following five phases to search for and identify relevant studies.
@@ 4.3. Search Items
To perform the manual search and establish QGS, we select four top-tier SE conferences (ICSE, ESEC/FSE, ASE, ISSTA) and two journals (TOSEM and TSE) and search for relevant papers involving both program repair and LLMs. We manually inspect and identify 25 papers from 2020 to 2024 that meet our criteria, which then form the basis for constructing QGS. Following prior surveys (Wang et al., 2022; Yang et al., 2022; Zhang et al., 2023b), we divide the search items used for searching papers automatically into two groups: (1) an APR-related group containing some commonly-used keywords related to program repair; and (2) an LLM-related group containing some popular keywords related to large language models or pre-trained models. Besides, as the APR community benefits from well-maintained forums and real-time updated works, we further refine our search items by concluding some frequent keywords from three sources: a community-driven website [^2] http://program-repair.org/bibliography.html, a living review of APR by Monperrus (Monperrus, 2022) and the most recent learning-based APR survey (Zhang et al., 2023b). Finally, we identify a search string that includes several LLM-related and APR-related keywords frequently appearing in program repair studies that use LLMs. The complete set of search keywords is shown in Table LABEL:tab:search_items.
Table 1. Search items used for retrieving LLM-based APR papers.
@@ 4.4.2. Quality Assessment
To ensure the inclusion of high-quality papers, we conduct a full-text inspection of the remaining papers to assess their relevance and quality within the LLM-based APR field. This step is adopted to mitigate the impact of substantial quality variations observed in the collected studies, particularly among very recent and non-peer-reviewed papers, and is consistent with common practices in most recent LLM-based SE surveys (Zhang et al., 2023c, 2025a). Following Wang et al. (Wang et al., 2022), we design ten quality assessment questions, where each paper is independently rated by two authors on a three-tier scale:“yes” (1 point), “partial” (0.5 points), or “no” (0 points). Papers scoring below the threshold of 8 points are excluded from further analysis, resulting in 167 papers being retained. Any disagreement is resolved through discussion, and when necessary, by consultation with the third author. The designed quality assessment criteria (QAC) are presented in Table 3. It is worth noting that, for QAC4, given the emerging nature of this research area and the rapid evolution of LLMs, many recent works have not yet completed the peer-review process. Therefore, we also consider papers from arXiv that satisfy our quality assessment criteria. Before inclusion, we carefully check whether each arXiv paper had been subsequently published by comparing its title, authorship, and content. If a published version is found, only the peer-reviewed version is retained; otherwise, the arXiv version is included and explicitly marked as such in our dataset. This strategy aligns with the common practice in recent LLM-based SE surveys, and ensures our work remains both comprehensive and up to date while maintaining a high standard of quality.
@@ 4.5. Snowballing Search
To ensure maximal comprehensiveness of the collected papers, we additionally employ the snowballing search  (Watson et al., 2022) to manually incorporate any potential papers that are overlooked during our previous search process. The snowballing search is a common practice to examine the references (i.e., backward snowballing) or citations (i.e., forward snowballing) within a list of papers to discover additional relevant works. Specifically, we scrutinize every reference and citation within the collected papers to assess their relevance to our research objectives. This meticulous manual analysis of all potential papers through the entire study selection process ultimately leads to the inclusion of 189 papers in our survey. Through this process, we identified 21 additional studies, leading to the inclusion of 189 final papers in our survey. All included papers and their bibliographic information are publicly available in our replication repository for transparency and reproducibility (Author, s).
@@ 5.4. What are the types of publication contributions?
We categorize collected papers according to their main contributions into four categories: new technique or methodology, empirical study, benchmark, and human study, as illustrated in Table 5. It is worth noting that, although most studies are assigned to a single category, a small number of studies make equally substantial contributions to multiple categories. We find that 118 relevant papers are published, with the aim of proposing a novel repair approach or framework with LLMs to address various issues in the APR community. Besides, 56 papers concentrate on conducting empirical studies to explore the actual benefits of LLMs in fixing various bugs, such as the potential of fine-tuning LLMs in vulnerability repair (Zhang et al., 2023d; Ge and Wu, 2023). We further notice 15 relevant studies constructing new benchmarks to evaluate the performance of LLMs and three papers (Ge and Wu, 2023; Meem et al., 2024) administering a survey to offer insights into how practitioners or developers think about and employ LLMs to fix software bugs in practice.
@@ 6.2. What approaches are employed to optimize LLMs for program repair?
2025 6 (Ehsani et al., 2025b), (Liu et al., 2025), (Orvalho et al., 2025), (Nong et al., 2025), (Lin et al., 2025), (Zhang et al., 2025d)
@@
2025 17 (Deligiannis et al., 2023), (Garg et al., 2023), (Huang et al., 2025c), (Mündler et al., 2024), (Charalambous et al., 2023), (Xu et al., 2023), (Xia et al., 2025a), (Yuan et al., 2024), (Zhang et al., 2024c), (Steenhoek et al., 2025), (Cao et al., 2023), (Hu et al., 2025), (Kong et al., 2024), (Li et al., 2025), (Ouyang et al., 2025), (Parasaram et al., 2025), (Kang et al., 2023)
@@ 6.2.4. Agent
The second category involves repository-level repair agents, which target real-world software issues, guided by the development of new benchmarks, such as SWE-bench (Jimenez et al., 2024). Unlike traditional approaches (Zhang et al., 2023b) that often rely on predefined conditions (e.g., assuming known bug locations, failure-triggering tests, or single-function repair assumptions), these agents require end-to-end bug repair across full repositories, requiring reasoning across multiple files, dependencies, and test suites. Recent studies have proposed a series of SWE-oriented techniques to address the challenges of repository-level repair, including designing specialized agent-computer interfaces (e.g., SWE-Agent (Yang et al., 2024a)), encoding the whole repository via AST representations (e.g., AutoCodeRover (Zhang et al., 2024d)), enabling multi-agent collaboration (e.g., MAGIS (Tao et al., 2024)), and developing generalist and specialist agents (e.g., OpenHands (Wang et al., 2024b)). More recently, unlike most prior agent-based approaches that rely on offline-trained LLMs and fixed prompting strategies, the agent-based paradigm opens new opportunities for incorporating online learning into APR systems (Mu et al., 2025; Xia et al., 2025b) , enabling agents to continuously refine their behavior during the repair process in evolving software projects.
@@ 7. RQ3: What repair scenarios have been facilitated by LLMs?
2025 22 (Bouzenia et al., 2024) (Cao et al., 2023) (Ehsani et al., 2025a) (Feng et al., 2025) (Hu et al., 2025) (Kong et al., 2024) (Kong et al., 2025) (Lee et al., 2024) (Li et al., 2025) (Lin et al., 2025) (Luo et al., 2024) (Macháček et al., 2025) (Ouyang et al., 2025) (Parasaram et al., 2025) (Ruiz et al., 2025) (Xu et al., 2025) (Yang et al., 2024e) (Ye et al., 2025) (Huang et al., 2025b) (Zhang et al., 2025d) (Kang et al., 2023) (Silva et al., 2023)
@@
2025 11 (Ehsani et al., 2025b) (Guo et al., 2025) (Ma et al., 2025a) (Mündler et al., 2024) (Antoniades et al., 2024) (Ouyang et al., 2024b) (Wei et al., 2025) (Xia et al., 2025a) (Tao et al., 2024) (Wang et al., 2024b) (Ma et al., 2025b)
@@ 7.1. Semantic Bugs
2025 ReinFix (Zhang et al., 2025d) ICSE GPT-3.5,GPT-4o,GPT-4 Java URL
@@ 7.4. Repository-level Issues
2025 Agentless (Xia et al., 2025a) FSE/ESEC Approach GPT-4o Python URL
@@
Repository-level issues refer to real-world issues (such as bug reports or feature requests) in large-scale repositories. Recent LLM-based APR research increasingly targets repository-level issues, whose resolution requires long-context understanding over a large codebase, accurate fault localization from informal reports, dependency, and code edits across multiple modules, and robust validation of the resulting pull request. A prominent benchmark for this scenario is SWE-bench (Jimenez et al., 2024), which comprises 2294 real GitHub issues with paired ground-truth pull requests from 12 popular Python repositories. In addition, several other datasets, such as SWT-bench (Mündler et al., 2024), OmniGIRL (Guo et al., 2025), and SWE-bench M (Yang et al., 2024b), are used to evaluate LLMs under varying APR settings. To address repository-level repair challenges, researchers have explored two main categories of approaches: procedural (Xia et al., 2025a; Ouyang et al., 2024b) and agentic (Zhang et al., 2024d; Antoniades et al., 2024; Chen et al., 2024b; Wang et al., 2024b). In particular, similar to few-shot or zero-shot prompting approaches, procedural approaches rely on expert-designed workflows where LLMs run at a fixed pipeline, and every choice is hard-coded. Agentic approaches query LLMs to autonomously plan and execute each choice by equipping LLMs with various external tools